In a recent poll, 230 people were asked if they liked dogs, and 80% said they did. Find the margin of error of this poll, at the 99% confidence level.

Give your answer to three decimals
Answer #1

Solution :

Given that,

n = 230

Point estimate = sample proportion = \hat p = 0.80

1 - \hat p = 0.20

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.80 * 0.20) / 230)

= 0.068

The margin of error of this poll, at the 99% confidence level is 0.068.